OpenStudy (anonymous):

2-5m=13 steps again...

OpenStudy (mertsj):

Subtract 2 from both sides first.

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

11

OpenStudy (mertsj):

Now divide both sides by -5

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

but that comes out to 2.2... and it needs to come out as a whole number..

OpenStudy (mertsj):

That would be - 2.2 and it cannot come out as a whole number unless you copied the problem wrong.

OpenStudy (anonymous):

okay.. the real question is.. 13=5m-2 i didnt think it would natter if you flopped it around..

OpenStudy (mertsj):

You didn't "flop it around". You changed 5m-2 to its opposite. That's like saying, ok, I'll just change this 10 to -10

OpenStudy (anonymous):

it still comes out 2.2...

OpenStudy (mertsj):

Add 2 to both sides.

OpenStudy (mertsj):

\[13+2=5m-2+2\]

OpenStudy (mertsj):

\[15=5m\]

OpenStudy (mertsj):

Divide both sides by 5

OpenStudy (anonymous):

and the answers 3?

OpenStudy (mertsj):

\[\frac{15}{5}=\frac{5m}{5}\]

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

okay tell me if i get this right to this next question.. -15=2t-11

OpenStudy (anonymous):

so would you add 15 to both sides?

OpenStudy (mertsj):

ok

OpenStudy (mertsj):

Remember you are trying to get the variable by itself. The variable has been multiplied by 2 and then had 11 subtracted from it. How would you "undo" those things so that the variable can be by itself?

OpenStudy (anonymous):

so.. then would you do -2t then? so then you could get rid of it?

OpenStudy (mertsj):

You "undo" the stuff in reverse order. The variable was multiplied by 2 and 11 was subtracted. To undo that add 11 and then divide by 2

OpenStudy (anonymous):

oh.. i see that now. sorry.. this is really frustrating..

OpenStudy (mertsj):

Well I know it is but if you keep at it, all of a sudden the light will come on.

OpenStudy (anonymous):

yeah, after i figure out what to either use to "undo" then i got the rest of the problem.. its just i never know what number to either subtract from or add.. you know what i mean?

OpenStudy (mertsj):

I do know what you mean. Just always focus on getting the variable by itself by "undoing" the stuff. What did you get for this problem?

OpenStudy (anonymous):

i got, t=-2

OpenStudy (mertsj):

that was very good. Do you have another one?

OpenStudy (anonymous):

yes i do, 7k-5=-19

OpenStudy (mertsj):

Ok. What do you think comes first?

OpenStudy (anonymous):

and this is the part i mean i have troubles with.. like i never know if you would add -5bor -19..

OpenStudy (mertsj):

What is the variable?

OpenStudy (mertsj):

I don't mean the solution. I just want you to identify the variable. Is it x or y or ...what is it?

OpenStudy (anonymous):

the variable is k

OpenStudy (mertsj):

Yes. Good. Now look at it. Is it multiplied or divided by something?

OpenStudy (anonymous):

7

OpenStudy (mertsj):

Exactly. So at some point we're going to have to reverse that by dividing by 7.

OpenStudy (mertsj):

But first, was anything added or subtracted to or from the k?

OpenStudy (anonymous):

yeah because the problem is 7k-5

OpenStudy (mertsj):

Yes you are correct but I want you to tell me specifically what the number is and whether it was added or subtracted.

OpenStudy (mertsj):